The length of a rectangle is 5m longer than the width. What are the dimensions if the area is 36m^2?
Answer by JulietG(1812)

L * W = Area
You're given both the length and the area. L=W+5
36 m^2 = (W+5) * W
Easiest way to solve this is to find the factors of 36
36,1
18,2
9,4
6,6
Which of those pairs fits the equation L = W+5? L is the first factor, W is the second.